1) A charge of +2uC is placed 4m away from a charge of -3uC. Determine the magnitude and direction of the force of the charges


1
Expert's answer
2021-10-25T15:24:59-0400

We can find the electrostatic force between two charges from the Coulomb's Law:


"F=\dfrac{kq_1q_2}{r^2},""F=\dfrac{9.0\times10^9\ \dfrac{N\times m^2}{C^2}\times2\times10^{-6}\ C\times(-3\times10^{-6}\ C)}{(4\ m)^2},""F=-3.4\times10^{-3}\ N."

The magnitude of the electrostatic force between two charges is "3.4\times10^{-3}\ N". Since the two charges are opposite in sign they will attract each other, therefore, the direction of the electrostatic force is attractive (also, the sign minus indicates that the direction is attractive).